Circular No. 8521

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005bx IN MCG +12-13-19 D. Rich, Hampden, ME, reports his discovery of an apparent supernova

(mag ∼ 17.8) on unfiltered CCD frames taken with a 0.31-m reflector on Apr. 27.10 and 30.08 UT. The new object is located at α = 13^{h}50^{m}24^{s}.95, δ = +68^{o}33′19′′.4 (equinox 2000.0), which is ∼ 1′′.5 west and 2′′.4 north of the center of MCG +12-13-19. Nothing is visible at this location on CCD frames taken by Rich on Feb. 28.20 and Apr. 11.08 (limiting mag ∼ 18.6) or on Palomar Digital Sky Survey images from 1993 and 1996 (limiting red and blue magnitudes fainter than 19.0).

SUPERNOVA 2005bf IN MCG +00-27-5 L. Wang, Lawrence Berkeley National Laboratory; and D. Baade, Eu-

ropean Southern Observatory (ESO), on behalf of the collaboration on supernova polarimetry program, report that the type-Ic supernova 2005bf (IAUC 8507, 8520) was observed at the ESO’s Very Large Telescope (+ FORS1; range 323–840 nm; resolution 0.26 nm/pixel) in polarimetry mode on Apr. 30. Preliminary inspection of the spectra indicates a transition to a type II-or type-Ib supernova. Assuming a recession velocity for the host galaxy of 5670 km/s (Falco et al. 1999, PASP 111, 438), conspicuous lines from He I 501.6-, 587.6-, 667.8-, and 706.5-nm are observed at velocities around 6200 km/s. The observations also show strong lines of Ca II H and K and the Ca II infrared triplet at velocities around 10000 km/s. An absorption line is detected at 624.8 nm — which, if identified with the Si II 635.5-nm line, gives an expansion velocity of ∼ 5000 km/s; the same line, however, can be identified as Hα at a velocity of 14400 km/s. Similarly, absorption features are observed at 413.9 and 463.4 nm, which can be identified with Hγ and Hβ, respectively, at a velocity of 14000 km/s. A blend of Fe II 516.0-nm and Fe III 515.6-nm is also present in the spectra, at velocity around 5200 km/s. The supernova shows polarized spectral features in excess of 2 percent across some spectral lines. Observations at all wavelengths are strongly encouraged.

GRB 050502 D. T. Durig, University Of The South, measures GRB 050502 (cf. GCN

3322, 3325, 3327) at α = 13^{h}29^{m}46^{s}.31, δ = +42^{o}40′27′′.2 (equinox 2000.0); approximate CCD V magnitudes: May 2.179 UT, 19.8; 2.182, 20.8; 2.186, 20.3; 2.197, 20.9; 2.201, 21.4; 2.208, 21.0; 2.264, 21.5; 2.264, 22.0.

2005 May 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green